OpenStudy (anonymous):

What is the geometric mean of 3 and 21?

OpenStudy (anonymous):

sqrt(3*21) = sqrt(9*7) = 3*sqrt(7)

OpenStudy (anonymous):

Ok I understand! Thank you!